# Attention Retrieves, MLP Memorizes: Disentangling Trainable Components in the Transformer

### v2

@@ 5 Discussion
As the attention matrix in MixiT is static and input-independent, MixiT cannot adapt to each input and form specific circuits such as induction heads [37]. Induction head circuits look for recent occurrences of the current token, and attends to the tokens that follow with increased probability. This allows the standard transformer to easily adapt to the input context in language modeling. Hence, it is no surprise that MixiT lags behind standard transformer for language modeling. It is perhaps more surprising that the perplexity comes close to that of the standard transformer.
The near-perfect performance on certain algorithmic tasks in Table 1 suggests that induction heads and other specialized input-dependent circuits are not required on these tasks, hence providing an axis for task separation.
Note that MixiT is able to perform well on the Yelp reviews dataset, despite the complexity of language used in reviews. This supports the hypothesis that sentiment can be largely judged by the collective token embeddings, as opposed to next-token prediction in language modeling tasks, which requires retrieving specific details from the context, a task induction heads are apt at.
@@
Interestingly, as shown in in Figure 3 and demonstrated by its performance, the Frozen-QK model can solve the retrieval task by forming induction heads.
@@
Previous works have highlighted the importance of MLPs in storing knowledge [13, 21, 20, 19, 55, 11, 31, 33]. They posit that specific facts from the training data are stored in specific knowledge neurons. These works support our findings, in that MLPs are crucial in knowledge memorization. However, our work does not prescribe knowledge localization, i.e. we don’t attribute memorizing specific facts to specific neurons.
Our work adds characterization on knowledge memorization in more recent works on knowledge circuits [54] and query localization [10], where MLPs and attention are found to collaborate on knowledge memorization, e.g. where attention selects the appropriate knowledge neurons depending on the query. Our work shows that even with static random attention weights, such as in Frozen-QK, attention and MLPs can still collaborate effectively, as evident in language modeling perplexities similar to that of the standard transformer, and the formation of specific input-dependent circuits. However, our results show that the role of MLPs in memorization outweighs that of attention, as evident by the fact that Frozen-MLP achieves much worse accuracy than Frozen-QK or MixiT (4).
@@
Zhong and Andreas [59] studies the random transformer, wherein they train only the embedding and unembedding layers, and leave the intermediate layers fixed at random initialization. The random transformer was found to be able to perform nontrivial algorithmic tasks. There are several notable differences between MixiT and the random transformer. The principal difference is that the random attention matrix $I+W^{M}_{\ell}-\bar{W}^{M}_{\ell}$ in MixiT is input-independent, and hence cannot develop specialized circuits that adapt to the input, whereas such circuits are possible for the random transformer. Moreover, the MLPs in MixiT are trained, whereas they remain frozen in the random transformer. These differences allow us to study the roles different trained components play in the transformer architecture.
Furthermore, through in depth hyperparameter searches, we find that the random transformer does not scale well with respect to depth, confirming some of the original findings [59]. However, MixiT, with its carefully normalized random attention matrix designed to preserve signal propagation, does scale with respect to depth, suggesting that the random transformer suffers from signal propagation challenges and rank collapse without appropriate shaping [14, 34, 35].
@@
Notably, some of these results strengthen the argument for empirical approaches for architecture design found in previous work [39, 8]. In particular, Poli et al. [39] uses performances in various synthetic tasks to design powerful hybrid architectures. Our results suggest that in certain circumstances some architectural components are entirely responsible for some basic operations (e.g. information retrieval) while in some other there is a more coordinated effort. Thus, overall this favours hybrid architectures with diverse basic components.
@@ 6 Related work
Several studies explore simplified Transformers with frozen or random components. Notably, random Transformers with fixed layers but trainable embeddings can solve many algorithmic tasks [59]. Similarly, replacing attention with fixed matrices - as in Synthesizer [46] or FNet [28] — retains competitive performance on certain benchmarks, suggesting learned attention is not always necessary. However, retrieval tasks often require flexible, input-dependent attention to form induction-like circuits. While their attention is also random, unlike our MixiT model it is input-dependent and so does not isolate the specific tasks for which attention is and is not needed. We relate our results to their work in detail in Section 5. Other past work has studied properties of other randomly frozen or lightly trained models, e.g. convolutional networks [25, 43, 3], largely without focusing on specific tasks.
@@
While there is a long line of work studying signal propagation in deep neural networks [44, 40, 27, 53], it was only more recently that Martens et al. [30] introduced the concept that modifying activation functions can significantly improve stability of signal propagation, and leading to rapid training of large scale vision models without using normalization and skip connections [57]. This was later understood to yield a stable scaling limit, which is characterized by an SDE of the covariance matrix [29]. The covariance SDE framework is then used to understand how to design and shape general non-linearities like the Transformer self-attention [35], resolving the rank collapse issue [14, 34] and has shown strong performance despite a much simplified Transformer block [23]. Our theoretical result Theorem 2.1 also follows from this framework.
@@
Several works have investigated the capabilities of Transformers in the classes of tasks analyzed here, including arithmetic [36]. The role of feedforward layers in memorization in the Transformer architecture has been studied in Geva et al. [18] and their inductive biases and scaling properties have been scrutinized [5]. In this context, our work shows their relevance in conjuction with trainable or fixed attention. Orthogonal to our work, memorization has also been studied to understand generalization in neural networks [56, 4, 2].
@@
Other closely related work is in the mechanistic interpretability, which aims to understand LLMs via examining and modifying their internals [6]. Closely related is work related to identifying and understanding the behavior of induction heads [37, 16, 7] and in-context learning [9]. Our work demonstrates that the performance separation between input-dependent and input-independent attention is largely driven by the latter’s inability to form induction heads. Finally, Meng et al. [31] which reverse-engineer how different Transformer components support behaviors like memorization, retrieval, and generalization.
@@
A last area that has seen significant effort at understanding attention is that of efficient Transformers [47]. While we study what happens when attention is replaced with a fixed input-independent matrix, this field has studied various useful aspects of the attention matrix such as attention sinks [51] and compression [26]. As our work demonstrates that for many tasks the full power of input-dependent attention is not needed, it may have its own implications for efficiency, e.g. by removing the need for the KV-cache.
@@ 7 Conclusion
In this work, we isolated the roles of the architecture’s two trainable components by selectively freezing parameters and introducing MixiT, whose attention maps are fixed and random. Fixed attention suffices for many algorithmic tasks. However, adaptive attention is essential for associative recall and language modelling, as MixiT fails on long-range retrieval and yields higher perplexity. However, memorization resides in the MLPs, with attention contributing significantly less. Together, these results support a heterogeneity hypothesis: MLPs and learnable attention mechanism are required for the full behavioural range of modern large language models.
Limitations and future work
Our work studied basic algorithmic, mathematical reasoning, and language modeling tasks. It is an exciting future direction to extend our study to more complex tasks such as human reasoning, which demands a wide-ranging skillset that does not fall neatly on one side of the retrieval-memorization spectrum. We used relatively small-scale models; extending the analysis to larger networks and datasets will test whether the observed separation of concerns persists at a larger scale. Future work might study hybrid training schedules in which only a subset of architectural modules remain trainable – or are gradually unfrozen – which may strike an even better accuracy–efficiency trade-off. Finally, probing how these architectural choices interact with emerging interpretability and safety techniques constitutes an exciting avenue for further research.
@@ Acknowledgments
Appendix A Proof of Theorem 2.1
@@
Consider the MixiT recursion $h_{\ell+1}=\text{Attn}(h_{\ell})$ defined by Equation 2.3 at initialization. Then as the width $n$ and depth $d$ go to infinity with $\frac{d}{n}\to\bar{\tau}>0$, the upper triangular entries of the covariance matrix $\Phi_{\ell}=\frac{1}{n}h_{\ell}^{\top}h_{\ell}$ flattened to a vector in $\mathbb{R}^{m(m+1)/2}$ converges to the solution of the following SDE
@@
To complete the proof, we will invoke the Markov chain convergence to SDE results in the Skorohod topology, see for example Li et al. [29, Proposition A.6], which gives us the desired result.
Appendix B Additional Experimental Details
B.1 Hyperparameter selection
We select for the optimal hyperparameters for each task using a grid search. The search ranges for each task are as follows, determined a priori depending on task complexity, e.g. language modeling is inherently more complex than memorization.
@@ B.1 Hyperparameter selection
We use a sequence length of 256 for Yelp sentiment classification and language modeling tasks. Language modeling tasks are trained for 40,000 steps.
For the $k$-hop induction heads task, following [42], we generate data using a sequence length of 100, and a maximum of $k=16$ hops. We use a training set of 100,000 samples, a test set of 100 samples, and train for 5,000 steps.
@@
Appendix C Further Discussion

### v3

@@ 5 Discussion
As the attention matrix in MixiT is static and input-independent, MixiT cannot adapt to each input and form specific circuits such as induction heads (Olsson et al., 2022). Induction head circuits look for recent occurrences of the current token, and attends to the tokens that follow with increased probability. This allows the standard transformer to easily adapt to the input context in language modeling. Hence, it is no surprise that MixiT lags behind standard transformer for language modeling. It is perhaps more surprising that the perplexity comes close to that of the standard transformer.
The near-perfect performance on certain algorithmic tasks in Table 3 suggests that induction heads and other specialized input-dependent circuits are not required on these tasks. Hence the MixiT performance on a given task can serve as a litmus test for whether in-context reasoning is required for that task. For instance, MixiT is able to perform well on the Yelp reviews dataset, despite the complexity of language used in reviews. This indicates that sentiment can be largely judged by the collective token embeddings, as opposed to next-token prediction in language modeling tasks, which requires retrieving specific details from the context, a task induction heads are apt at.
@@
Interestingly, as shown in Figure 3 and demonstrated by its performance shown in Table 1 and Table 2, the Frozen-QK model can solve the retrieval task by forming induction heads.
@@
These results naturally raises the question: how expressive is Frozen-QK? To answer this, we prove the following result showing Frozen-QK can approximate a wide class of functions:
Theorem 5.1 (Universal Approximation of Frozen-QK).
Every continuous causal function with compact support can be approximated arbitrarily well by one layer of multihead attention and MLP, where query and key weight matrices are frozen at random initialization.
The proof can be found in Appendix B. In summary, the proof leverages the fact that standard transformers are universal approximators of such functions (Yun et al., 2020), and lifts this universal approximation to random feature transformers. Note that in practice, there are also the value weights and the MLP layers in the attention module, so we do expect the representation power to be even stronger, which is evidenced by empirical results.
Furthermore, the proof sheds light on why MixiT cannot be a universal approximator: as $Q_{k}$ and $K_{k}$ are input-independent in MixiT, the random feature $g_{i}=x\ \text{Softmax}(Q_{k}^{\top}K_{k})$ in MixiT is linear. Hence non-linear functions in $x$, such as the induction head function, cannot be approximated.
@@
Previous works have highlighted the importance of MLPs in storing knowledge (Dai et al., 2022; Geva et al., 2023, 2022, 2021; Yu and Ananiadou, 2024; Chughtai et al., 2024; Meng et al., 2022; Merullo et al., 2024). They posit that specific facts from the training data are stored in specific knowledge neurons. These works support our findings, in that MLPs are crucial in knowledge memorization. However, our work does not prescribe knowledge localization, i.e. we don’t attribute memorizing specific facts to specific neurons.
Our work adds characterization on knowledge memorization in more recent works on knowledge circuits (Yao et al., 2024) and query localization (Chen et al., 2025), where MLPs and attention are found to collaborate on knowledge memorization, e.g. where attention selects the appropriate knowledge neurons depending on the query. Our work shows that even with static random attention weights, such as in Frozen-QK, attention and MLPs can still collaborate effectively, as evident in language modeling perplexities similar to that of the standard transformer, and the formation of specific input-dependent circuits. However, our results show that the role of MLPs in memorization outweighs that of attention, as evident by the fact that Frozen-MLP achieves much worse accuracy than Frozen-QK or MixiT (4).
@@
Zhong and Andreas (2024) studies the random transformer, wherein they train only the embedding and unembedding layers, and leave the intermediate layers fixed at random initialization. The random transformer was found to be able to perform nontrivial algorithmic tasks. There are several notable differences between MixiT and the random transformer. The principal difference is that the random attention matrix $I+W^{M}_{\ell}-\bar{W}^{M}_{\ell}$ in MixiT is input-independent, and hence cannot develop specialized circuits that adapt to the input, whereas such circuits are possible for the random transformer. Moreover, the MLPs in MixiT are trained, whereas they remain frozen in the random transformer. These differences allow us to study the roles different trained components play in the transformer architecture.
Furthermore, through in depth hyperparameter searches, we find that the random transformer does not scale well with respect to depth, confirming some of the original findings (Zhong and Andreas, 2024). However, MixiT, with its carefully initialized random attention matrix designed to preserve signal propagation, does scale with respect to depth, suggesting that the random transformer suffers from signal propagation challenges and rank collapse without appropriate shaping (Dong et al., 2021; Noci et al., 2022, 2023).
@@
These observations provide an explanation for the subspace selection hypothesis in (Zhong and Andreas, 2024), namely rank collapse with respect to model depth. Rank collapse refers to the phenomenon where the representations of different inputs become more and more similar as model depth increases, and has previously been studied in e.g. (Dong et al., 2021; Noci et al., 2022, 2023). To substantiate this, we analyze the covariance between hidden representations for the language modeling task on Wikitext. This covariance is calculated between the last layer hidden representations within a sequence (higher covariance means the representations of different tokens are more similar), then averaged across sequences.
Model \ # Layer 2 4 8 16 32
Random Transformer .00218 .01450 .02075 .01946 .05438
MixiT .00088 .00260 .00210 .00154 .00104
Table 7: Covariance between hidden representations of different tokens for models of different depths.
Table 7 shows that the covariance for the Random Transformer steadily increases as the model size increases in depth, but remains steady for MixiT. This directly implies representation collapse in the Random Transformer, where the representation of different tokens become increasingly similar as the model size grows in depth.
This covariance degeneracy is especially catastrophic for language modeling, where deep models are required to achieve good performance. This helps explain why the Random Transformer struggles with language modeling and solving tasks when the model size increases beyond a certain number of layers (Zhong and Andreas, 2024).
@@
Notably, some of these results strengthen the argument for empirical approaches for architecture design found in previous work Poli et al. (2024); Carstensen et al. (2025). In particular, Poli et al. (2024) uses performances in various synthetic tasks to design powerful hybrid architectures. Our results suggest that in certain circumstances specific architectural components are entirely responsible for some basic operations (e.g. information retrieval), while in others it is a more coordinated effort (e.g. memorization).
@@ 6 Related work
Several studies explore simplified Transformers with frozen or random components. Notably, random Transformers with fixed layers but trainable embeddings can solve many algorithmic tasks Zhong and Andreas (2024). Similarly, replacing attention with fixed matrices - as in Synthesizer Tay et al. (2021) or FNet Lee-Thorp et al. (2021) — retains competitive performance on certain benchmarks, suggesting learned attention is not always necessary. However, retrieval tasks often require flexible, input-dependent attention to form induction-like circuits. While their attention is also random, unlike our MixiT model it is input-dependent and so does not isolate the specific tasks for which attention is and is not needed. We relate our results to their work in detail in Section 5. Other past work has studied properties of other randomly frozen or lightly trained models, e.g. convolutional networks (Jarrett et al., 2009; Saxe et al., 2011; Arora et al., 2019), largely without focusing on specific tasks. In addition, (Hassid et al., 2022) also studies static attention, but uses the learned attention matrices over a reference corpus, and hence is not completely data-free, and does not have a stable signal propagation guarantee as MixiT. Indeed the completely data-free attention variants (Hassid et al., 2022) tested, without any reference corpus, perform poorly. Our work shows that such a variant’s performance depends heavily on the task: on whether specialized circuits need to be formed.
@@
While there is a long line of work studying signal propagation in deep neural networks Schoenholz et al. (2016); Poole et al. (2016); Lee et al. (2017); Yang and Schoenholz (2017), it was only more recently that Martens et al. (2021) introduced the concept that modifying activation functions can significantly improve stability of signal propagation, and leading to rapid training of large scale vision models without using normalization and skip connections Zhang et al. (2022). This was later understood to yield a stable scaling limit, which is characterized by an SDE of the covariance matrix Li et al. (2022). The covariance SDE framework is then used to understand how to design and shape general non-linearities like the Transformer self-attention Noci et al. (2023), resolving the rank collapse issue Dong et al. (2021); Noci et al. (2022) and has shown strong performance despite a much simplified Transformer block He and Hofmann (2023). Our theoretical result Theorem˜2.1 also follows from this framework.
@@
Several works have investigated the capabilities of Transformers in the classes of tasks analyzed here, including arithmetic Nogueira et al. (2021). The role of feedforward layers in memorization in the Transformer architecture has been studied in Geva et al. (2020) and their inductive biases and scaling properties have been scrutinized Bachmann et al. (2023). In this context, our work shows their relevance in conjuction with trainable or fixed attention. Orthogonal to our work, memorization has also been studied to understand generalization in neural networks Zhang et al. (2016); Arpit et al. (2017); Anagnostidis et al. (2022).
@@
Other closely related work is in the mechanistic interpretability, which aims to understand LLMs via examining and modifying their internals (Bereska and Gavves, 2024). Closely related is work related to identifying and understanding the behavior of induction heads (Olsson et al., 2022; Edelman et al., 2024; Bietti et al., 2023) and in-context learning Chan et al. (2022). Our work demonstrates that the performance separation between input-dependent and input-independent attention is largely driven by the latter’s inability to form induction heads. (Chen et al., 2024) also studies different components of the transformer model, but from a model dynamics point of view, analyzing differences in gradients between components, and on synthetic tasks such as indirect object identification and factual recall. Finally, Meng et al. (2022) which reverse-engineer how different Transformer components support behaviors like memorization, retrieval, and generalization.
@@
A last area that has seen significant effort at understanding attention is that of efficient Transformers (Tay et al., 2022). While we study what happens when attention is replaced with a fixed input-independent matrix, this field has studied various useful aspects of the attention matrix such as attention sinks (Xiao et al., 2024) and compression (Kim et al., 2024). As our work demonstrates that for many tasks the full power of input-dependent attention is not needed, it may have its own implications for efficiency, e.g. by removing the need for the KV-cache.
@@ 7 Conclusion
In this work, we designed a spectrum of model architectures to systematically study the components of a transformer model. We found that, surprisingly, trainable attention is not required to form specialized circuits, and Frozen-QK can indeed perform well on language modeling. We also identified an obstruction towards scaling depth-wise in prior random models, and designed a principled remedy. Our work sheds important light on the functionalities of different model components, such as attention and MLPs, and shows how each is crucial for different tasks.
Limitations and future work.
Our work studied basic algorithmic, mathematical reasoning, and language modeling tasks. It is an exciting future direction to extend our study to more complex tasks such as human reasoning, which demands a wide-ranging skillset that does not fall neatly on one side of the retrieval-memorization spectrum. Future work might study hybrid training schedules, in which only a subset of architectural modules remain trainable – or are gradually unfrozen – which may strike an even better accuracy–efficiency trade-off. Finally, probing how these architectural choices interact with emerging interpretability and safety techniques constitutes an exciting avenue for further research.
@@ Acknowledgments
Appendix A Proof of Theorem˜2.1
@@
Consider the MixiT recursion $h_{\ell+1}=\text{Attn}(h_{\ell})$ defined by Equation˜2.3 at initialization. Then as the width $n$ and depth $d$ go to infinity with $\frac{d}{n}\to\bar{\tau}>0$, the upper triangular entries of the covariance matrix $\Phi_{\ell}=\frac{1}{n}h_{\ell}^{\top}h_{\ell}$ flattened to a vector in $\mathbb{R}^{m(m+1)/2}$ converges to the solution of the following SDE
@@
To complete the proof, we will invoke the Markov chain convergence to SDE results in the Skorohod topology, see for example Li et al. [2022, Proposition A.6], which gives us the desired result.
Appendix B Proof of Theorem 5.1
Theorem (Universal Approximation of Frozen-QK).
Every continuous causal function with compact support can be approximated arbitrarily well by one layer of multihead attention and MLP, where query and key weight matrices are frozen at random initialization.
Proof.
The proof will follow the universal approximation theory for random feature regression in Banach spaces by Neufeld and Schmocker [2023], where we will set up the relevant Banach spaces.
Let the space of input sequences of length $m$, where each element is a vector in $V\subset\mathbb{R}^{n}$, where $V$ is compact, be denoted by $U=V^{m}$. The Banach space of continuous functions that map an input sequence from a compact set $K\subset U$ to an output sequence in $U$ is denoted by $(C(K,U),||\cdot||_{\infty})$, i.e. equipped with the sup-norm. The subspace of continuous causal functions, which we can denote as $C_{\text{causal}}(K,U)$, consists of all functions $f\in C(K,U)$ that satisfy the following property: For any given position $i\in\{1,\dots,m\}$ and any two input sequences $u=(u_{1},\dots,u_{m})$ and $v=(v_{1},\dots,v_{m})$ in the domain $K$:
$\text{If }u_{j}=v_{j}\text{ for all }j\leq i,\text{ then }(f(u))_{i}=(f(v))_{i}$
This condition ensures that the output at position $i$ only depends on the input up to position $i$. Since this subspace is a closed linear subspace of the Banach space $C(K,U)$, it is itself a Banach space with the same supremum norm.
We define a random feature model based on a single multi-head attention layer with causal masking. Let the input be a sequence of hidden states $h\in\mathbb{R}^{n\times m}$, where $m$ is the sequence length and $n$ is the embedding dimension. The layer has $N_{h}$ attention heads. Conceptually, each attention head $i\in\{1,\dots,N_{h}\}$ generates a single, matrix-valued random feature. This feature is the product of the hidden space $h$ and an attention pattern $\mathcal{A}_{i}$, which is a function that maps the input sequence $h$ to an $m\times m$ matrix. The randomness for each feature $\mathcal{A}_{i}$ comes from a pair of frozen weight matrices, $(W^{Q}_{i},W^{K}_{i})$, where $W^{Q}_{i},W^{K}_{i}\in\mathbb{R}^{n\times d_{k}}$. These are drawn independently for each head from a random distribution (e.g. $\mathcal{N}(0,1)$) at initialization and are not trained.
The random feature (attention pattern) for head $i$ is defined as:
$g_{i}=h\,\mathcal{A}_{i}(h;W^{Q}_{i},W^{K}_{i})=h\,\text{Softmax}\left(\frac{(W^{Q}_{i}h)^{\top}(W^{K}_{i}h)}{\sqrt{d_{k}}}+M_{\text{causal}}\right)\,$, (B.1)
where $M_{\text{causal}}$ is the causal mask matrix that prevents attention to future positions.
By Theorem 3 in Yun et al. [2020], Transformers are universal approximators for compactly supported sequence-to-sequence functions. Furthermore, as Neufeld and Schmocker [2023] shows that universal approximation can be lifted from deterministic feature functions to random feature functions in Banach spaces [Neufeld and Schmocker, 2023, Theorem 3.2], we can lift this universal approximation to the random features $g_{i}$, to approximate any causal function $f\in C_{\text{causal}}(K,U)$ to arbitrary precision.
Note that, given that both the value weight matrix $W^{V}_{i}$ of each head as well as the weights of the MLP layer are trainable, this is strictly more expressive than the random feature model we just constructed with $g_{i}$. Therefore one layer of self-attention and MLP must also be a universal approximator of causal functions, as desired.
Appendix C Additional Experimental Details
C.1 Hyperparameter selection
We select for the optimal hyperparameters, shown in Table 9, for each task using a grid search. The search ranges for each task are shown in Table 8, determined a priori depending on task complexity, e.g. language modeling is inherently more complex than memorization.
@@ C.1 Hyperparameter selection
Furthermore, we use a sequence length of 256 for Yelp sentiment classification and language modeling tasks. Language modeling tasks are trained for 40,000 steps.
For the $k$-hop induction heads task, following [Sanford et al., 2024], we generate data using a sequence length of 100, and a maximum of $k=16$ hops. We use a training set of 100,000 samples, a test set of 100 samples, and train for 5,000 steps.
For a fair comparison across models on the memorization task (Table 4), we use the same hidden dimension, with $L=2$, and $n_{h}=4$ heads. We train for $10,000$ steps with a learning rate of $0.005$. Similarly, for a fair comparison on the retrieval task, reported in Figure 2, we fix the embedding dimension at $n=1024$, with $4$ heads and $2$ layers, which suffice to learn induction heads [Elhage et al., 2021, Edelman et al., 2024]. We search for the optimal learning rate for each architecture. Note that these two sets of results are meant to capture model performance when the accuracies are not saturated at 100%, to give a meaningful comparison between models.
@@
Appendix D Further Discussion